# Collaborative and parametric insurance on the Ethereum blockchain

### v1

@@ Acknowledgements
Pierre-O’s work is conducted as part of the Research Chair DIALOG [^12] See https://chaire-dialog.fr/en/welcome/, under the auspices of the Risk Foundation, an initiative by CNP Assurances. His research is also supported by the ANR project BLOCKFI [^13] See https://anr.fr/Project-ANR-24-CE38-7885. Stéphane acknowledges support from the Research Chair ACTIONS [^14] See https://chaireactions.fr/ (funded by BNP Paribas Cardif), the Initiative de Recherche Sustainable Actuarial Science and Climatic Risk [^15] See https://sites.google.com/view/stephaneloisel/recherche/projets/actuariat-durable-et-risques-climatiques (funded by Milliman Paris), and the ANR project DREAMES [^16] See https://anr.fr/Projet-ANR-21-CE46-0002.

### v2

@@ Acknowledgements
Pierre-Olivier Goffard’s work is conducted as part of the Research Chair DIALOG [^10] See https://chaire-dialog.fr/en/welcome/, under the auspices of the Risk Foundation, an initiative by CNP Assurances. His research is also supported by the ANR project BLOCKFI [^11] See https://anr.fr/Project-ANR-24-CE38-7885. Stéphane Loisel acknowledges support from the Research Chair ACTIONS [^12] See https://chaireactions.fr/ (funded by BNP Paribas Cardif), the research initiative Sustainable Actuarial Science and Climatic Risks [^13] See https://sites.google.com/view/stephaneloisel/recherche/projets/actuariat-durable-et-risques-climatiques (funded by Milliman Paris), and the ANR project DREAMES [^14] See https://anr.fr/Projet-ANR-21-CE46-0002.